A 1575 kg car is pulling a trailer of a mass 982kg. If the car and trailer are accelerating at 1.07m/s^2 west, what force is the road exerting on the wheels of the car? What force does the trailer hitch exert on the trailer?
Answer #1

Solution :

Given :

mass of car (M) = 1575 kg

mass of trailer (m) = 982 kg

acceleration (a) = 1.07 m/s2

what force is the road exerting on the wheels of the car?

Fby road on wheels of car = ( m + M ) a = ( 1575 kg + 982 kg )(1.07 m/s2) = 2735.99 N

What force does the trailer hitch exert on the trailer?

Fby trailer hitch on the trailer = m a = (982 kg)(1.07 m/s2) = 1050.74 N